OpenStudy (anonymous):

3, -13, and 5 + 4i f(x) = x4 - 8x3 - 12x2 + 400x - 1599 f(x) = x4 - 200x2 + 800x - 1599 f(x) = x4 - 98x2 + 800x - 1599 f(x) = x4 - 8x3 + 12x2 - 400x + 1599

OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form.

ganeshie8 (ganeshie8):

lets do this in unconventional way

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

5+4i is one root so ?

OpenStudy (anonymous):

5-4i is also one

ganeshie8 (ganeshie8):

list all the roots once

OpenStudy (anonymous):

3, -13, 5-4i and 5+4i

ganeshie8 (ganeshie8):

\(\alpha + \beta + \gamma + \psi = \frac{-b}{a} \)

ganeshie8 (ganeshie8):

sum of roots equals that. Now just dont get confuse yet, just add all the roots u listed, and tell me wat u get

OpenStudy (anonymous):

f(x) = x4 - 200x2 + 800x - 1599

ganeshie8 (ganeshie8):

I want you add all the roots you listed and see wat u get

ganeshie8 (ganeshie8):

sum of roots : 3 + -13 + 5-4i + 5+4i = ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

-10

ganeshie8 (ganeshie8):

check again

OpenStudy (anonymous):

ahhhh

OpenStudy (dan815):

>_>

OpenStudy (dan815):

ZERO

ganeshie8 (ganeshie8):

yess

ganeshie8 (ganeshie8):

so 0 = -b/a, => b = 0

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

there will not be x^3 term in the polynomial.

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

So, answer is either B or C

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

lets figure out the x^2 coefficient

ganeshie8 (ganeshie8):

Ive told you this is unconventional way, so if u dont get it, its okay :)

OpenStudy (anonymous):

ok, how do you do that

OpenStudy (dan815):

why does this work all the time?

ganeshie8 (ganeshie8):

\(\alpha \beta + \beta \gamma + \gamma \psi +\psi \alpha = \frac{c}{a}\)

OpenStudy (dan815):